OpenStudy (anonymous):

tan(sin^-1x + cos^-1y) Answer in terms of x and y

OpenStudy (anonymous):

\[\tan (\sin^{-1}x +\cos^{-1} y)\]

OpenStudy (gorv):

\[\sin ^{-1}x+\cos^{-1}x=\frac{ \pi }{ 2 }\]

OpenStudy (gorv):

replace sin